# leftmargin=0pt, rightmargin=0pt, backgroundcolor=bggray, middlelinecolor=black, roundcorner=3 *[subfigure]position=bottom

### v1

@@ B.5. Suggestions and Toxicity, Sentiment, and Character Agency
[…]
Agency, toxicity, and sentiment ratings in stories about characters of various genders. We see few differences between stories written with and without suggestions for each gender.
@@
toxicity $t(337)=-0.672$ $0.6982$
sentiment - $t(72)=-0.066$ $0.9634$
Student sentiment $t(337)=2.516$ $0.0462$
@@
Beyond the attributes we designed the predictive text model to suggest to the user, we also consider off-the-shelf classification of stories toxicity [^10] https://huggingface.co/martin-ha/toxic-comment-model, sentiment (Camacho-collados et al., 2022), and character agency (Wan et al., 2023). Here, we consider writing scenarios where the character of interest’s gender is pre-specified before the participants begin writing. We compare between classifier output for stories where characters are masc-coded vs fem-coded when participants are or are not provided with suggestions (Figure 16 and Table 17).
@@
We find no significant differences in sentiment between suggestion conditions in the Teacher and Town Hall scenarios regardless of suggestions and in the Student scenario without suggestions. However, in the Student scenario, we see significantly lower sentiment ratings for stories written about “John” than “Abby” when the stories are written with suggestions ($t(337)=2.516$, $p_{\text{FDR}}\approx 0.0462$). This means that the predictive text model may have had a bias towards suggesting more positive continuations about Abby than John or that participants were more likely to accept positive suggestions about Abby. Regardless of mechanism, we see that, in this scenario, predictive text suggestions widened the gap in sentiment between genders.
@@
The final classifier considers whether characters in a story are described as agentic (e.g., being a natural leader) vs communal (e.g., being a well-liked member of a group) (Wan et al., 2023). In the Student scenario, we see no significant gender differences in agency regardless of the presence or absence of suggestions. In the Town Hall scenario, we see significantly higher agency in stories about “Thomas” than “Rebecca” when they are written with suggestions ($t(332)=3.852$, $p_{\text{FDR}}\approx 0.0010$). We see a similar trend in the Teacher scenario. Here the increased agency for “Mr. Brown” is significant with suggestions ($t(333)=3.44$, $p_{\text{FDR}}\approx 0.0040$) and marginally significant without suggestions ($t(71)=2.424$, $p_{\text{FDR}}\approx 0.0625$). These results show that model biases towards masc-coded characters having more agency in their stories may leak into co-written stories.
Appendix C Annotation and Validation
@@ C.1. Annotation Prompts and Instructions
[…]
Example annotation prompt. The system prompt is “You will be given a premise and a hypothesis. Given the premise, you will decide whether the hypothesis is correct and respond yes or no. If there is not enough information, respond no. Return only the word yes or no, and nothing else.” We then include a couple in context examples of the annotation task, and finally the participant’s (partial) story and a hypothesis to annotate.
@@
Teacher likeable According to the speaker, {Mrs./Mr.} Brown has a likeable personality.
@@ C.2. Common Contributing Words
As discussed in subsection 5.2, we annotate at the word level to determine which words (either included in the story or proposed and rejected by the model) contributed to the gender, likeability, confidence, etc of story characters. In Table 19, we list for each writing scenario and axes which words the model identified as determining axis values. Note that the same word may appear for both values of an axis. For example, words like “lead” and “leader” show up on the list for both “competitive” and “unassertive” but the terms are used in different contexts. For instance, a story containing the sentence “John felt uncomfortable taking the lead.” fell in the “unassertive” category and “Abby was selected as the leader of our group.” fell in the opposite.
@@
Teacher likeable favorite (103); kind (25); patient (25); best (24); most (22); inspired (18); favourite (15); inspiration (10)
@@ C.3. Human Evaluation Details and Instructions
To validate the LLM annotations of the human or co-written stories, we collect human annotations from $10$ annotators. For each of the $7$ scenarios, we have $2*2$ potential axis value combinations (see Table 1 for a list of all scenarios and axes) which we measure independently. For each of these measurements (e.g. the character “Mr. Brown” has a likeable personality), the value can true or false/unspecified. This leaves us $7*4*2=56$ unique measurement values made about the set of stories. We collect $560$ random sets of these unique story measurement values. Each annotator is asked to annotate $56$ stories for single axis values, but these tasks are randomized between annotators to avoid them learning patterns about how many “true” and “false” values there should be per scenario, axis, etc. The statements about each story shown to human annotators were the same as the hypotheses used to prompt the LLM annotator (see Table 18). We include the instructions to human annotators below:
@@
Please mark the statement as false if it is either untrue or is unspecified using your best judgement about what can be “reasonably” inferred from the story. For example, for a story where the character’s likeable vs repellent personality is not explored at all, please mark “False”. For a story where the character is seen by the narrator as likeable or is shown to be likeable in one anecdote, one could argue that you cannot infer whether they are inherently likeable overall, but these should still be marked as “True”.
Appendix D User Study Interface
@@
[…]
What is your level of English proficiency? What is your age? What is your gender identity? Please provide your Prolific ID. Do you have any feedback about the study?
@@
[…]
Your response has been recorded. Thank you for participating in the study.
Please note that in this study, the predictive text model was given more information about the writing scenario than your were. For example, the model may have been told to make a given character ”confident” or ”unconfident”. Some of these scenarios reflect potentially harmful stereotypes (e.g., that female characters are less confident than male characters) which should not be taken as fact.
You will soon be redirected to Prolific. Please wait…

### v2

@@ B.5. Suggestions and Toxicity, Sentiment, and Character Agency
Bar charts showing agency, toxicity, and sentiment ratings in stories about characters of different genders. Differences between stories written with and without suggestions are minimal.
@@
toxicity t(337)=0.672t(337)=-0.672 $0.6982$
sentiment - t(72)=0.066t(72)=-0.066 $0.9634$
Student sentiment $t(337)=2.516$ $0.0462$ *
@@
Beyond the attributes the predictive text model was controled to suggest, we also consider off-the-shelf classification of stories toxicity [^16] https://huggingface.co/martin-ha/toxic-comment-model, sentiment (camacho-collados-etal-2022-tweetnlp), and character agency (wan-etal-2023-kelly). Here, we consider writing scenarios where the character of interest’s gender is cued in the story prefix. We compare between classifier output for stories where characters are masc-coded vs fem-coded when participants are or are not provided with suggestions (Figure 17 and Table 17).
@@
We find no significant differences in sentiment between suggestion conditions in the Teacher and Town Hall scenarios regardless of suggestions and in the Student scenario without suggestions. However, in the Student scenario, we see significantly lower sentiment ratings for stories written about “John” than “Abby” when the stories are written with suggestions ($t(337)=2.516$, $p_{\text{FDR}}\approx 0.0462$, $d=0.273$). This means that the predictive text model may have had a bias towards suggesting more positive continuations about Abby than John or that participants were more likely to accept positive suggestions about Abby. Regardless of mechanism, we see that, in this scenario, predictive text suggestions widened the gap in sentiment between genders.
@@
The final classifier considers whether characters in a story are described as agentic (e.g., being a natural leader) vs communal (e.g., being a well-liked member of a group) (wan-etal-2023-kelly). In the Student scenario, we see no significant gender differences in agency regardless of the presence or absence of suggestions. In the Town Hall scenario, we see significantly higher agency in stories about “Thomas” than “Rebecca” when they are written with suggestions ($t(332)=3.852$, $p_{\text{FDR}}\approx 0.0010$, $d=0.422$). We see a similar trend in the Teacher scenario. Here the increased agency for “Mr. Brown” is significant with suggestions ($t(333)=3.44$, $p_{\text{FDR}}\approx 0.0040$, $d=0.377$) and marginally significant without suggestions ($t(71)=2.424$, $p_{\text{FDR}}\approx 0.0625$, $d=0.568$). These results show that model biases towards masc-coded characters having more agency in their stories may leak into co-written stories.
B.6. Prefix Similarity on Attribute Selection
The uniform information density (UID) hypothesis states that people prefer to uniformly distribute information throughout language production when possible to maintain the same message (Frank and Jaeger, 2008; Jaeger and Levy, 2006; meister-etal-2021-revisiting). An implication of UID is that low-probability words may be more likely followed by high-probability words (and vice versa). In our scenario, this could potentially impact the selection of pro-stereotypical (likely higher probability) versus anti-stereotypical (likely lower probability) as a function of the probability of the word (or phrase) that came before. Under this interpretation, participants are not (only) choosing a pro-stereotypical word because it is higher probability, but because the preceding word is low probability. For example, in spoken language, fillers are a common way to add additional time before a low probability event, such as “Bill married his [uh] long-time boyfriend” vs “Bill married his long-time girlfriend.”
The UID interpretation in our setting suggests that anti-stereotypical suggestions may not be taken because the user was not planning on the low probability event (“boyfriend”) and by the time the anti-stereotypical suggestion arrived (after “long-time”) it was too late to make an appropriate high-probability selection in advance of the low-probability continuation. Then, in this interpretation, the user is even more strongly guided to select the high-probability continuation that they had in mind (“girlfriend”), irrespective of any anti-stereotypical suggestion.
While this interpretation is possible—and could be an interesting avenue for future research—we expect that its effect is rather small, for two reasons. First, measured UID effects tend to be quite small. For example, UID effects on log likelihood are on the order of, at most, $\pm 0.15$ nats in meister-etal-2021-revisiting, in comparison to probability differences on the order of as much as $\pm 20$. Second, in our study, the story prefix is minimally edited within scenarios (e.g., changing only “Mr. Brown” vs “Mrs. Brown”). This means that early in each story, the information density between conditions should be roughly the same. Thus, any tokens that appear very early in the story will, by definition, have nearly the same prefix and therefore nearly the same past information density.
(a) Gender
(b) ABC Traits
Figure 18. What number writing action determines the given attribute in each story? The first writing action (after the “start” action) is highlighted in orange.
Histograms showing the story position of writing actions that determine each attribute in co-written stories. For gender, a spike occurs at position 1, indicating gender is often specified in the first new token. For ABC traits, contributions are distributed further along in the stories.
More specifically, for gender (18(a)), we have that about $35\%$ of gender-defining tokens are written or accepted in the first action (counting the “start” action as action 0) with $63\%$ of gender-defining tokens coming from the first five actions. For the Detective scenario, we see a large portion of gender-defining tokens as the second action where participants write a name like “Detective X” instead of just “X”. For the Wedding scenario, we see some slightly longer phrases before the partner’s gender is written out from common phrases like “his highschool sweetheart X”, “her soulmate X”, etc. Overall, we see that a large portion of gender-defining tokens are written with nearly identical contexts (within writing scenarios), leading the prior information density to be roughly equal when the gender decision is made.
For attributes beyond gender (18(b)), we see more variability in when the ABC trait is specified. For most scenarios, the ABC trait cannot easily be specified in the first token. In the Teacher scenario, the teacher’s likability is never determined by the first written or accepted token. Instead, in about $31\%$ of stories, participants specify the teacher’s likability on the second token with phrases like “my favorite”, “the worst”, etc. While this means likability in this case was often determined with prefixes with similar information density, for the remaining scenarios, ABC traits may have been determined with prefixes that have more variable information density which may have influenced participant behavior.
Overall, we see that there are some attributes writing scenarios that were largely determined early enough in stories that information density is very consistent between conditions. We argue that our other findings cannot be explained through effects of uniform information density alone.
Appendix C Annotation, Validation, and Additional Participant Details
@@ C.1. Annotation Prompts and Instructions
Example annotation prompt showing the system instruction: participants receive a premise and hypothesis, and decide whether the hypothesis is correct by responding “yes” or “no.” The figure includes in-context examples and the participant’s partial story with a hypothesis to annotate.
@@
Teacher likable According to the speaker, {Mrs./Mr.} Brown has a likable personality.
@@ C.2. Common Contributing Words
As discussed in subsection 5.2, we annotate at the word level to determine which words (either included in the story or proposed and rejected by the model) contributed to the gender, likability, confidence, etc of story characters. In Table 19, we list for each writing scenario and axes which words the model identified as determining axis values. Note that the same word may appear for both values of an axis. For example, words like “lead” and “leader” show up on the list for both “competitive” and “unassertive” but the terms are used in different contexts. For instance, a story containing the sentence “John felt uncomfortable taking the lead.” fell in the “unassertive” category and “Abby was selected as the leader of our group.” fell in the opposite.
@@
Teacher likable favorite (103); kind (25); patient (25); best (24); most (22); inspired (18); favourite (15); inspiration (10)
@@ C.3. Human Evaluation Details and Instructions
To validate the LLM annotations of the human or co-written stories, we collect human annotations from $10$ annotators. For each of the $7$ scenarios, we have 222*2 potential axis value combinations (See Table 1 for a list of all scenarios and axes) which we measure independently. For each of these measurements (e.g. the character “Mr. Brown” has a likable personality), the value can true or false/unspecified. This leaves us 742=567*4*2=56 unique measurement values made about the set of stories. We collect $560$ random sets of these unique story measurement values. Each annotator is asked to annotate $56$ stories for single axis values, but these tasks are randomized between annotators to avoid them learning patterns about how many “true” and “false” values there should be per scenario, axis, etc. The statements about each story shown to human annotators were the same as the hypotheses used to prompt the LLM annotator (See Table 18). We include the instructions to human annotators below:
@@
Please mark the statement as false if it is either untrue or is unspecified using your best judgement about what can be “reasonably” inferred from the story. For example, for a story where the character’s likable vs repellent personality is not explored at all, please mark “False”. For a story where the character is seen by the narrator as likable or is shown to be likable in one anecdote, one could argue that you cannot infer whether they are inherently likable overall, but these should still be marked as “True”.
C.4. Additional Participant Details
As we discussed in subsection 4.3, we do not restrict study participation based on country to allow for a more diverse set of English proficiency levels. We show a breakdown of participant nationality in Figure 20.
Figure 20. Self-reported nationality of participants included in the final analysis (i.e., those who completed the study and passed the attention check). Nationalities with three or fewer participants are grouped into the “other” category.
Bar chart showing the nationality of participants in decreasing order: South Africa: 72; United Kingdom: 59; United States: 49; Poland: 42; Hungary: 22; Portugal: 21; Mexico: 18; Canada: 16; Nigeria: 14; Italy: 13; Chile: 10; Spain: 8; Greece: 7; Australia: 5; Zimbabwe: 4; Czech Republic: 3; Brazil: 3; other: 39
Appendix D Additional Model Details
As we discuss in section 5, we use Llama models to generate both predictive text suggestions and to annotate the resulting stories. We provide suggestion generation code at https://github.com/ctbaumler/predictive_text and summary details below:
  • Model: Llama 2-Chat 7B (Touvron et al., 2023)
  • Quantization: None
  • Sampling / decoding: Greedy; select top-3 candidates from raw logits (temperature = 0). Each suggestion is decoded to the next word boundary.
  • Prompt format: See Figure 4 for an example and https://github.com/ctbaumler/predictive_text for the full prompt set.
The details for the annotator model are as follows:
  • Model: Llama 3 70B (AI@Meta, 2024) (instruction-tuned variant)
  • Quantization: None
  • Scoring method: We compute the softmax over the model’s logits for the tokens “yes” and “no,” and use the resulting probabilities $p(\text{yes})$ and $p(\text{no})$ for analysis. No sampling is involved. For story-level annotations, we apply a cutoff of $0.8$ to determine whether an attribute is present. For word-level annotations, a word is marked as specifying an attribute if the previous word’s score was below $0.8$, the current word’s score is above $0.8$, and the difference between the two exceeds $0.3$. (See subsection 5.2 for details and an example.)
  • Prompt format: See Figure 19 for a single example and https://github.com/ctbaumler/predictive_text for the full set of prompts.
Appendix E User Study Interface
@@
Screen asking “What is your level of English proficiency? What is your age? What is your gender identity? Please provide your Prolific ID. Do you have any feedback about the study?”
@@
Screen showing “Your response has been recorded. Thank you for participating in the study. Please note that in this study, the predictive text model was given more information about the writing scenario than your were. For example, the model may have been told to make a given character ”confident” or ”unconfident”. Some of these scenarios reflect potentially harmful stereotypes (e.g., that female characters are less confident than male characters) which should not be taken as fact. You will soon be redirected to Prolific. Please wait…”